OpenStudy (anonymous):

. Solve using perfect square factoring patterns. y2 + 14y + 49 = 0 {–7} {0, –1} {7} {7, –7}

OpenStudy (anonymous):

why is y^2 + 14y + 49 a perfect square trinomial?

OpenStudy (skullpatrol):

The factoring pattern you are trying to match is $$\huge a^2+2ab+b^2$$

OpenStudy (anonymous):

perfect square factoring patterns looks like this: a^2 + 2ab + b^2 = (a+b)^2 you have: y2 + 14y + 49 do you see how the places of each term are like the perfect square places?

OpenStudy (skullpatrol):

$$\huge a^2+2ab+b^2=(a+b)^2$$ $$\huge y^2 + 14y + 49$$ $$\huge y^2 + 2(7)y + 7^2=(?+?)^2$$